A Carnot engine has a power output of 139 kW. The engine operates between two reservoirs at 16◦C and 658◦C.

(A)How much thermal energy is absorbed each hour? Answer in units of J

(B)How much thermal energy is lost per hour? Answer in units of J.


1
Expert's answer
2022-05-02T13:43:36-0400

"\frac{Q_2}{Q_1}=\frac{T_2}{T_1},"

"Q_1=448~J,"

"Q_3=Q_2-Q_2=309~J."